OpenStudy (anonymous):

i can not understand why -1 raised to the power 8 is -1

OpenStudy (cggurumanjunath):

1

OpenStudy (cggurumanjunath):

http://prntscr.com/2m53ja

OpenStudy (cggurumanjunath):

it is +1 go through above link.

OpenStudy (mathmale):

(-1)^8 = ( (-1)^2 )^4= (1)^4 = 1